# Invariance of deep image quality metrics to affine transformations

### v2

@@ II-D Sensitivities of metrics and humans for the different distortions
In Figure 7, the human thresholds indicate that the human sensitivity order (indicated by the vertical lines) is: scale, translation, rotation, RG, and YB illuminants. However, the metric shown returns the following order (given by the slopes of the curves): scale, rotation, translation, YB, and RG illuminants. Even if the ordering doesn’t strictly match, it is to be noted that this metric maintains more sensitivity to geometric than chromatic distortions.
@@
III Experimental Settings
In this section, we review the selected affine transformations, models, and databases that we will use in the experiments. Code available to test other metrics publicly [^2] https://github.com/Rietta5/InvarianceTestIQA
III-A Affine Transformations
@@
Some examples of affine transformations that follow Equation 7 are geometric contraction, expansion, dilation, reflection, rotation, or shear; and some examples following Equation 8 are changes in contrast, luminance, and illuminant. In this work, we are going to focus specifically on translation, rotation scale, and illuminant changes. For each tested affine transformation, the original images are modified in the following ways:
@@
  • Illuminant changes: We have desaturated the original images and modified the illuminant for 20 hues in 8 saturations, i.e. angle with respect to the $x$ axis, and distance with respect to the central white point respectively. Following the distribution at the locus as indicated in Figure 8.
@@
Figure 9: Examples of affine transformations applied to the original images. From left to right it shows the original image, translation (150 pixels right), rotation (15 degrees), scale (factor of 0.8), and illuminant change (hue 17, purple, with maximum intensity).
III-B Datasets
@@
Figure 10: Example of an original image and with the mosaic effect to fill the image when affine transformations modify it out of the original dimensions. Once the mosaic is created, the affine transformation is applied and cropped, as indicated by the red frame, recovering the initial dimensions. An example of the final result can be seen in Figure 9.
@@
III-C Metrics
We choose different metrics because of their relevance in the context of perceptual metrics, either because of their age and widespread use, or because of the good results obtained in other works and applied to other areas of study. In this work, the following will be used:
@@ III-C Metrics
IV Results
Table II summarizes the numerical results of the experiments. For each geometric affine transformation, we present the discrimination thresholds per metric, $\theta_{\tau}^{M}$. In the case of illuminant changes, we show the errors made with respect to the interpolated MacAdam and the experimentally fitted ellipses. As the ellipses are very similar (Figure 5), the errors are very similar. In those where they did not coincide completely, the mean of the errors has been calculated and marked with an asterisk (*) in the Table. The Figures from which these numerical results are derived are given in Appendix B and Appendix C to avoid cluttering the main text.
As we are considering both the human thresholds extracted from the literature, $\theta_{\tau}^{H_{L}}$, and the threshold obtained by ourselves with natural images, $\theta_{\tau}^{H_{N}}$, a result is marked in bold if $\theta_{\tau}^{M}$ falls within $\theta_{\tau}^{H_{L}}$ and is underlined if it falls within $\theta_{\tau}^{H_{N}}$. Note that a result can be marked both ways at the same time. A metric with many highlighted boxes will be a metric with a good performance from the point of view of human invariance.
TABLE II: Distortion range of the metrics for each geometric affine distortion. For geometric transformations, intervals that match the literature threshold are marked in bold, while those that match the thresholds measured during the work are underlined. For illuminant changes, metrics that return a smaller error with respect to the fitted ellipses are marked in the same way as in geometric ones.
@@ IV Results
translation MNIST 0.025$\pm$0.003 0.030$\pm$0.004 0.05$\pm$0.03 0.034$\pm$0.008 0.031$\pm$0.004 0.05$\pm$0.03 0.12
CIFAR-10 0.033$\pm$0.004 0.032$\pm$0.004 0.05$\pm$0.03 0.036$\pm$0.019 0.05$\pm$0.02 0.13$\pm$0.12
TID13 0.027$\pm$0.003 0.028$\pm$0.003 0.034$\pm$0.009 0.08$\pm$0.10 0.0029$\pm$0.003 0.06$\pm$0.04 0.23
ImageNet 0.036$\pm$0.010 0.033$\pm$0.004 0.05$\pm$0.04 0.13$\pm$0.15 0.036$\pm$0.010 0.15$\pm$0.13
Rotation MNIST 2$\pm$1 2.9$\pm$1.5 3.7$\pm$2 1.2$\pm$0.8 4.0$\pm$1.7 10.7$\pm$8.1 3
CIFAR-10 1.1$\pm$0.6 1.5$\pm$0.7 1.4$\pm$1.2 0.7$\pm$0.4 3.6$\pm$1.8 1.7$\pm$0.9
TID13 0.10$\pm$0.03 0.11$\pm$0.04 0.10$\pm$0.05 0.1$\pm$0.3 0.11$\pm$0.04 0.09$\pm$0.08 3.6
ImageNet 0.19$\pm$0.09 0.15$\pm$0.06 0.15$\pm$0.11 0.11$\pm$0.05 0.22$\pm$0.09 0.2$\pm$0.5
Scale MNIST 1.10$\pm$0.04 1.19$\pm$0.12 1.2$\pm$0.2 1.089$\pm$0.010 1.16$\pm$0.07 1.3$\pm$0.2 1.03
CIFAR-10 1.052$\pm$0.006 1.055$\pm$0.007 1.061$\pm$0.007 1.054$\pm$0.007 1.067$\pm$0.018 1.06$\pm$0.03
TID13 1.047$\pm$0.006 1.048$\pm$0.006 1.057$\pm$0.007 1.062$\pm$0.008 1.040$\pm$0.005 1.059$\pm$0.007 1.03
ImageNet 1.0080$\pm$0.0010 1.0074$\pm$0.0009 1.0086$\pm$0.0011 1.00813$\pm$0.00010 1.0082$\pm$0.0010 1.009$\pm$0.005
Illum MNIST 0.03$\pm$0.06* - - 0.016$\pm$0.018* - 0.05$\pm$0.05* MacAdam
CIFAR-10 0.007*$\pm$0.009* 0.03$\pm$0.05* 0.03$\pm$0.03* 0.005$\pm$0.013* 0.05$\pm$0.13* 0.06$\pm$0.06*
TID13 0.010$\pm$0.010* 0.02$\pm$0.02* 0.02$\pm$0.02 0.02$\pm$0.02 0.05$\pm$0.09* 0.005$\pm$0.012* Our ellipse
ImageNet 0.008$\pm$0.010 0.02$\pm$0.02* 0.02$\pm$0.02 0.02$\pm$0.02 0.05$\pm$0.09* 0.005$\pm$0.012*
The results from Table II show that following our strong invariance criteria, there is no clear winner. No model shows the required robustness to affine transformations nor type of stimuli.
The comparison of the order between the human thresholds and the sensitivities of the metrics can be seen in Table III, where even if no metric shows complete human behavior, most of them maintain the chromatic ordering.
TABLE III: Results associated with the least demanding test, on the order of the sensitivities and their comparison with the human order. The first row of the table shows the order of the humans, which does not change depending on the database. Displayed in two different colors: In purple, the order of the geometric transformations and, in pink, the order of the YB and RG illuminants.
@@
Human S$>$R$>$T$>$RG$>$YB S$>$R$>$T$>$RG$>$YB S$>$T$>$R$>$RG$>$YB S$>$T$>$R$>$RG$>$YB
SSIM R$>$S$>$T$>$RG$>$YB YB$>$T$>$S$>$RG$>$R YB$>$S$>$R$>$T$>$RG YB$>$T$>$S$>$RG$>$R
LPIPS R$>$RG$>$YB$>$S$>$T R$>$YB$>$RG$>$S$>$T YB$>$RG$>$R$>$S$>$T YB$>$R$>$RG$>$T$>$S
DISTS RG$>$S$>$YB$>$R$>$T R$>$RG$>$YB$>$S$>$T R$>$RG$>$YB$>$S$>$T R$>$S$>$RG$>$YB$>$T
PerceptNet S$>$T$>$R$>$RG$>$YB T$>$S$>$RG$>$R$>$YB S$>$R$>$T$>$YB$>$RG T$>$S$>$R$>$YB$>$RG
PIM RG$>$YB$>$R$>$S$>$T R$>$RG$>$YB$>$S$>$T RG$>$YB$>$R$>$S$>$T YB$>$RG$>$R$>$S$>$T
V Discussion
As we have seen in Table II, bearing in mind that no model is highlighted for all related transformations, the conclusion is that there is no metric that behaves, in terms of invariance, like the human being. We can dissect the behavior for the different affine transformation as follows:
@@ V Discussion
Getting into the metrics’ performance regarding the ordering of the transformations from Table III, PerceptNet is the only metric that almost maintains the ordering of being more sensitive to geometric over chromatic transformations. All the metrics are generally more sensitive to YB than RG, which matches human performance in energy definition. All in all, no metric can be said to have human-like behavior even with the least strict test.
In addition, a collateral result that can be obtained from our work is specific invisibility thresholds for each metric. That is thresholds in units of each metric that indicate variations for which the distance is imperceptible to the human eye. These values are shown in Table IV.
TABLE IV: Metric-specific invisibility thresholds. Indicates distance variations for which changes between two images are imperceptible to humans. For example, a value equal to or less than 3.09 in PIM implies that these two images are indistinguishable from the observer.
@@
VI Conclussion
@@
Regarding human thresholds, we used both classical and explicitly measured values of the thresholds obtained using the same kind of natural images employed in perceptual metrics. Whereas for metric thresholds we propose a methodology to assign invisibility regions for any particular metric, which can be expressed in units of the metric themselves or in physical units to facilitate direct comparison with human psychophysics results.
@@
Making the demanding comparison between human and metric thresholds for a range of state-of-the-art deep image quality metrics shows that none of the studied metrics (both deep as well as RSMSE and SSIM) succeeds under these criteria: they do not reproduce all the human thresholds for the different affine transformations, and they do not reproduce the order of sensitivities in humans. No metric is capable of reproducing the thresholds of invisibility and human sensibilities.
This means that tuning the models exclusively to predict quality ratings may disregard other properties of human vision for instance invariances or invisibility thresholds.
@@ Acknowledgment
This work was supported in part by MICIIN/FEDER/UE under Grant PID2020-118071GB-I00 and PDC2021-121522-C21, in part by Spanish MIU under Grant FPU21/02256, in part by Generalitat Valenciana under Projects GV/2021/074, CIPROM/2021/056 and CIAPOT/2021/9 and in part by valgrAI - GVA. Some computer resources were provided by Artemisa, funded by the European Union ERDF and Comunitat Valenciana as well as the technical support provided by the Instituto de Física Corpuscular, IFIC (CSIC-UV).
Appendix A.
@@
Appendix C.
Figures 16, 17, 15, 14 show the results for the all databases for the illuminant transformations showing the final ellipse in the spectral locus. The title of each subfigure specifies the errors made with respect to the MacAdam ellipse at that point in space (left) and with respect to the experimentally obtained ellipse (right).
@@
Appendix D.
Figures 18, 19,20, and 21 show the results for the all databases for the comparison of sensibilities. The vertical lines show the human thresholds for each transformation, the different curves, the behaviour of each model. The order of the vertical lines is compared with the value of the slope. A steep slope implies higher sensitivity.
Figure 18: Sensitivity test results for the MNIST dataset. Table III shows in detail the order of the sensitivities together with the human ordering, where it can be seen that no metric follows the order correctly. However, in all cases the relative order of illuminant changes is respected, i.e. they are more sensitive to the RG direction than to the YB direction.
Figure 19: Sensitivity test results for the CIFAR10 dataset. Table III shows in detail the order of the sensitivities together with the human ordering, where it can be seen that no metric follows the order correctly.
Figure 20: Sensitivity test results for the TID13 dataset. Table III shows in detail the order of the sensitivities together with the human ordering, where it can be seen that no metric follows the order correctly.
Figure 21: Sensitivity test results for the ImageNet dataset. Table III shows in detail the order of the sensitivities together with the human ordering, where it can be seen that no metric follows the order correctly.

### v3

@@ 2.4 Sensitivities of metrics and humans for the different distortions
In Figure 8, the human thresholds (vertical lines) indicate that the human sensitivity order is: scale, translation, rotation, RG, and YB illuminants, where a lower threshold implies a higher sensitivity. However, the metric shown returns the following order given by the slopes of the curves (slope is directly related to the sensitivity): scale, rotation, translation, YB, and RG illuminants. Even if the ordering doesn’t strictly match, it is to be noted that this metric maintains more sensitivity to geometric than chromatic distortions.
@@
3 Experimental Settings
In this section, we review the selected affine transformations, models, and databases that we will use in the experiments. Code available to test other metrics is available publicly [^1] https://github.com/Rietta5/InvarianceTestIQA.
3.1 Affine Transformations
@@
Some examples of affine transformations that follow Equation 7 are geometric contraction, expansion, dilation, reflection, rotation, or shear; and some examples following Equation 8 are changes in contrast, luminance, and illuminant. In this work, we are going to focus specifically on translation, rotation, scale, and illuminant changes. For each tested affine transformation, the original images are modified in the following ways:
@@
  • Illuminant changes: We have desaturated the original images and modified the illuminant for 20 hues in 8 saturations, i.e. angle with respect to the $x$ axis, and distance with respect to the central white point, respectively. Following the distribution at the chromatic diagram as indicated in Figure 9.
@@
Figure 10: Examples of affine transformations applied to the original images. From left to right it shows the original image, translation (100 pixels right), rotation (15 degrees), scale (factor of 1.3), and illuminant change (hue 17, purple, with maximum intensity).
3.2 Datasets
@@
Figure 11: Example of an original image and with the mosaic effect to fill the image when affine transformations modify it out of the original dimensions. Once the mosaic is created, the affine transformation is applied and cropped, as indicated by the red frame, recovering the initial dimensions. This is an example of how the mosaic behaves if we want to move the image to the right. Further examples of the final results can be seen in Figure 10.
@@
3.3 Metrics
We choose different metrics because of their relevance in the context of perceptual metrics, either because of their age and widespread use, or because of the good results obtained in other works and applied to other areas of study. These metrics are only chosen and analyzed to showcase our proposed methodology, we do not intend this work to be an in-depth analysis of state-of-the-art metrics for image quality. To this effect, the following metrics will be used:
@@ 3.3 Metrics
4 Results
Table 2 summarizes the numerical results of the experiments. For each geometric affine transformation, we present the discrimination thresholds per metric, $\theta_{\tau}^{M}$. In the case of illuminant changes, we show the errors made with respect to the interpolated MacAdam and the experimentally fitted ellipses. As the ellipses are very similar (Figure 6), the errors are very similar. In those where they did not coincide completely, the mean of the errors has been calculated and marked with an asterisk (*) in the Table. The Figures from which these numerical results are derived are given in Appendix B and Appendix C to avoid cluttering the main text.
As we are considering both the human thresholds extracted from the literature, $\theta_{\tau}^{H_{L}}$, and the threshold obtained by ourselves with natural images, $\theta_{\tau}^{H_{N}}$, a result is marked in bold if $\theta_{\tau}^{M}$ falls within $\theta_{\tau}^{H_{L}}$ and is underlined if it falls within $\theta_{\tau}^{H_{N}}$. For the ellipses, we check if the experimental and MacAdam ellipses fall within the uncertainty interval obtained from the psychometric functions. Note that a result can be marked both ways at the same time. A metric with many highlighted boxes will be a metric with a good performance from the point of view of human invariance.
Table 2: Distortion range of the metrics for each geometric affine distortion. For geometric transformations, intervals that match the literature threshold are marked in bold, while those that match the thresholds measured during the work are underlined. For illuminant changes, metrics that return a smaller error with respect to the fitted ellipses are marked in the same way as in geometric ones.
@@ 4 Results
[…]
The results from Table 2 show that following our strong invariance criteria, there is no clear winner. No artificial model shows the required robustness to affine transformations nor type of stimuli.
The comparison of the order between the human thresholds and the sensitivities of the metrics can be seen in Table 3, where even if no metric shows complete human behavior, most of them maintain chromatic order in contrast to geometric transformations.
Table 3: Results associated with the least demanding test, on the order of the sensitivities and their comparison with the human order. The first row of the table shows the order of the humans, which does not change depending on the database. Displayed in two different colors: In purple, the order of the geometric transformations and, in pink, the order of the YB and RG illuminants.
@@
Human S>>R>>T>>RG>>YB S>>R>>T>>RG>>YB S>>T>>R>>RG>>YB S>>T>>R>>RG>>YB
SSIM R>>S>>T>>RG>>YB YB>>T>>S>>RG>>R YB>>S>>R>>T>>RG YB>>T>>S>>RG>>R
LPIPS R>>RG>>YB>>S>>T R>>YB>>RG>>S>>T YB>>RG>>R>>S>>T YB>>R>>RG>>T>>S
DISTS RG>>S>>YB>>R>>T R>>RG>>YB>>S>>T R>>RG>>YB>>S>>T R>>S>>RG>>YB>>T
PerceptNet S>>T>>R>>RG>>YB T>>S>>RG>>R>>YB S>>R>>T>>YB>>RG T>>S>>R>>YB>>RG
PIM RG>>YB>>R>>S>>T R>>RG>>YB>>S>>T RG>>YB>>R>>S>>T YB>>RG>>R>>S>>T
5 Discussion
5.1 Analysis of the results
As we have seen in Table 2, bearing in mind that no model is highlighted for all related transformations, the conclusion is that there is no metric that behaves, in terms of invariance, like the human being. We can dissect the behavior for the different affine transformation as follows:
@@ 5.1 Analysis of the results
Getting into the metrics’ performance regarding the ordering of the transformations from Table 3, PerceptNet is the only metric that almost maintains the ordering of being more sensitive to geometric over chromatic transformations. All the metrics are generally more sensitive to RG than YB, which matches human performance in energy definition. All in all, no metric can be said to have human-like behavior even with the least strict test.
In addition, a collateral result that can be obtained from our work is specific invisibility thresholds for each metric. That is, thresholds in units of each metric that indicate variations for which the distance is imperceptible to the human eye. These values are shown in Table 4.
Table 4: Metric-specific invisibility thresholds. Indicates distance variations for which changes between two images are imperceptible to humans. For example, a value equal to or less than 3.09 in PIM implies that these two images are indistinguishable for the observer.
@@
5.2 Known limitations and future work
As this work relies heavily on experimental measurements, multiple experimental error sources influence the invisibility thresholds measured for the artificial models. Going back to Figure 1, we can identify three different sources of uncertainty. These uncertainties were excluded from the representation to avoid cluttering, but have to be taken into consideration: (1) Human visibility thresholds (black line in Figure 1), $\theta^{H}_{\tau}$, are measured with experimental psychometric function and have an associated experimental error. (2) Threshold in the internal distance representation (yellow line in Figure 1), $\mathcal{D}_{\tau}$, are obtained via psychometric function and, as such, have an experimental uncertainty. (3) Evaluations of different metrics (continuous curves in blue, red, and green in Figure 1) are obtained by applying the transduction function to the responses of the metric. This transduction function is obtained by performing a least squares fit to a given dataset (TID13), inducing another experimental error.
In order to mitigate the experimental uncertainties aforementioned, there are different avenues of work that could be taken into the future: (1, 2) increasing the number of individuals and image samples used to calculate the psychometric functions would reduce their experimental uncertainty. In particular, (2) would see more improvement because currently, there are few images in lower DMOS ranges. (3) Using a different database with more distortions and images would lead to a more precise least-squares fitting, thus reducing its uncertainty.
One could say that another one of the caveats of the methodology proposed in Section 2.2.1 is using TID13, as it does not contain affine transformations in its range of distortions. A clear path looking forward would be repeating the TID13 experiment but including these affine transformations to conform a bigger and more general dataset which could be used to repeat the methodology and re-measure the threshold in the common internal representation. It has to be said that this was not done in this paper because it would imply a lot of experimentation time that was not the main focus of the project. All in all, we consider this would make the proposed methodology more robust but we do not foresee a drastical change in the results.
6 Conclusion
@@
Regarding human thresholds, we used both classical and explicitly measured values of the thresholds obtained using the same kind of natural images employed in perceptual metrics. Whereas for metric thresholds, we propose a methodology to assign invisibility regions for any particular metric, which can be expressed in units of the metric themselves or in physical units to facilitate direct comparison with the results of human psychophysics.
@@
Making the demanding comparison between human and metric thresholds for a range of well-known and stablished deep image quality metrics shows that none of the studied metrics (both deep-learning-based as well as RSMSE and SSIM) succeeds under these criteria: they do not reproduce all the human thresholds for the different affine transformations, and they do not reproduce the order of sensitivities in humans. No metric is capable of reproducing the thresholds of invisibility and human sensibilities.
This means that tuning the models exclusively to predict quality ratings may disregard other properties of human vision as invariances or invisibility thresholds.
@@ Acknowledgment
This work was supported in part by MICIIN/FEDER/UE under Grant PID2020-118071GB-I00 and PDC2021-121522-C21, in part by Generalitat Valenciana under Projects GV/2021/ 074, CIPROM/2021/056, CIAPOT/2021/9, CIACIF/2023/223 and in part by valgrAI - GVA. Some computer resources were provided by Artemisa, funded by the European Union ERDF and Comunitat Valenciana as well as the technical support provided by the Instituto de Física Corpuscular, IFIC (CSIC-UV).
Appendix A Appendix A.
@@
figure*[!h]
Results of the different metrics in the experiments for the geometric affine transformations. On these graphs we will apply what we have seen in Figure 1 to create intervals to compare with the human thresholds. In the different columns, we have the results differentiated by database (and in the rows, by affine transformation (translation, rotation and scale).
Appendix C Appendix C.
Figures 18, 19, 17, 16 show the results for the all databases for the illuminant transformations showing the final ellipse in the chromatic diagram. The title of each subfigure specifies the errors made with respect to the MacAdam ellipse at that point in space (left) and with respect to the experimentally obtained ellipse (right).
@@
Appendix D Appendix D.
Figures 20, 21,22, and 23 show the results for the all databases for the comparison of sensibilities. The vertical lines show the human thresholds for each transformation, the different curves, the behaviour of each model. The order of the vertical lines is compared with the value of the slope. A steep slope implies higher sensitivity.
Figure 20: Sensitivity test results for the MNIST dataset. Table 3 shows in detail the order of the sensitivities together with the human ordering, where it can be seen that no metric follows the order correctly. However, in all cases the relative order of illuminant changes is respected, i.e. they are more sensitive to the RG direction than to the YB direction.
Figure 21: Sensitivity test results for the CIFAR10 dataset. Table 3 shows in detail the order of the sensitivities together with the human ordering, where it can be seen that no metric follows the order correctly.
Figure 22: Sensitivity test results for the TID13 dataset. Table 3 shows in detail the order of the sensitivities together with the human ordering, where it can be seen that no metric follows the order correctly.
Figure 23: Sensitivity test results for the ImageNet dataset. Table 3 shows in detail the order of the sensitivities together with the human ordering, where it can be seen that no metric follows the order correctly.
[…]